# PDFxTMDLib: A High-Performance C++ Library for Collinear and Transverse Momentum Dependent Parton Distribution Functions

### v3

@@ Abstract
Collinear parton distribution functions (cPDFs) and transverse momentum dependent distributions (TMDs) are essential for calculating cross sections in high-energy physics, particularly within collinear and $k_{t}$-factorization frameworks. Currently, there exists two libraries, such as LHAPDF and TMDLib, to obtain these physical objects. However, there are limitations in both libraries, especially for TMDs, such as restricted customization and extensibility. Users are limited to the implementations provided by these libraries and cannot easily support unconventional PDFs. Additionally, no standard TMD library currently supports calculations of QCD coupling and uncertainties, which are crucial for precise phenomenological studies.
@@ I Introduction
For TMDLib, these problems are more extensive. Currently, there is no accepted standard for TMD sets, specifically regarding the interpolation grid shape or a common info file format. This library is, in fact, a combination of different components, where each handles its specific TMD set under a common Abstract Programming Interface (API). Due to this lack of a standard procedure, supporting a new TMD set is a non-trivial task. Additionally, the absence of integrated options to calculate uncertainties or the QCD coupling makes working with this library challenging. Therefore, a clear need exists within the community for a new library to address these shortcomings.
@@ II.2.1 The PDFSet Class
  • alphaQCD(mu2), Computes the strong coupling $\alpha_{s}$ at a given scale $\mu^{2}$.
@@ II.5 Grid Data Formats and File Handling
PDFxTMDLib maintains compatibility with lhagrid1 file formats of LHAPDF while introducing a new file format for TMDs which is an extension of lhagrid1 format to address the lack of standardization in the TMD community. We name the new file format lhagrid_tmd1. This extended format maintains structural similarity to lhapdf1 but accommodates the additional dimensionality of TMDs by organizing data in grids of $x$, $k_{t}^{2}$, $\mu^{2}$, and flavors. For implementation simplicity, the current version utilizes a single subgrid structure, in contrast to the multiple subgrids supported in the lhapdf1 format.
The current implementation supports several TMD sets, including PB-LO-HERAI+II-2020-set1, PB-LO-HERAI+II-2020-set2, PB-NLO+QED-HERAI+II-set1, PB-NLO+QED-HERAI+II-set2, PB-NLO-HERAI+II-2023-set2-qs=0.74, and PB-NLO-HERAI+II-2023-set2-qs=1.04. The library’s support of TMD sets will expand in the future, with updates published at www.pdfxtmdlib.org. Additionally, an example of how to produce this file format is also available at the official website of this library.
The aforementioned TMD sets are generated by grid files of TMDLib library, where we have developed a C++ program that convert the TMD grids from their original file formats to the newer lhagrid_tmd1 file format. It should be mentioned that their original YAML meta data info file is also modified to support uncertainty and QCD coupling calculation. Importantly, PDFxTMDLib maintains also some supports for TMD sets of TMDLib with the specialized IReader, and IInterpolator implementations for allflavorUpdf format. However, it should be noted that due to lack of standard in TMDLib this support is partial.
@@
In this structure, $<nnnn>$ denotes a four-digit member identifier. For instance, ‘CT18NLO_0003.dat’ represents the third member of the ‘CT18NLO’ PDF set. The central metadata for each set is stored in a YAML-formatted info file bearing the set name.
@@ III.1 Drell-Yan Process Simulations with PDFxTMDLib and LHAPDF
To assess the validity of PDFxTMDLib, we compare the total and differential cross sections against those from LHAPDF. The total cross section for the Drell-Yan process was measured as $1.695\times 10^{-6}$ mb for both libraries, with statistical uncertainties of $1.453\times 10^{-10}$ mb, indicating identical results. Differential cross sections are also presented as histograms of the dilepton invariant mass, rapidity, and transverse momentum, exhibited excellent agreement between the two libraries. These results are presented in Figure 6, which overlays the distributions and confirms their consistency across all kinematic variables. This alignment shows that PDFxTMDLib is indeed a reliable alternative to LHAPDF for Drell-Yan simulations. In the next subsection, we keep a closer eye, on validity of PDFxTMDLib for both cPDFs and TMDs by comparing their distributions against the LHAPDF, and TMDLib.
Figure 6: Comparison of differential cross sections for the Drell-Yan process simulated with PYTHIA using PDFxTMDLib and LHAPDF with the MSHT20nlo_as120 PDF set. The distributions for the rapidity (top left), transverse momentum (top right), and invariant mass (bottom) of the dilepton system show excellent agreement between the two libraries.
@@ IV Installation and Usage Guide
This section provides detailed instructions for installing PDFxTMDLib and integrating it into research workflows. This section is one of the important sections for the users to help them to start to use this library. In addition of this section, if there were any further questions regarding building, and using this libary, users can ask questions via email, or official github address of the repository https://github.com/Raminkord92/PDFxTMD/discussions.
@@ IV.1.2 Build Process
mkdir build
cd build
cmake -DCMAKE_BUILD_TYPE=Release -S ..
cmake --build .
@@ IV.1.3 Installation
To install the library system-wide (recommended for Linux/macOS), execute:
cmake --install .
This installs headers, libraries, and CMake configuration files to standard system locations. On Linux/macOS, this typically requires administrative privileges.
@@ V Conclusion
Numerical validation, including Drell-Yan process simulations with PYTHIA and direct comparisons with LHAPDF and TMDLib, demonstrates that PDFxTMDLib delivers consistent and accurate results, alongside modest performance improvements in specific scenarios. The library further enhances phenomenological research by introducing novel features for TMDs, such as uncertainty quantification and QCD coupling calculations, which were previously unavailable in a standardized form. Additionally, we also introduce a new lhagrid_tmd1 file format for TMDs which is an extension of lhagrid file format to facilitate and standardize calculations for TMDs.

### v4

@@ Abstract
Collinear parton distribution functions (cPDFs) and transverse momentum dependent distributions (TMDs) are essential for calculating cross sections in high-energy physics, particularly within collinear and $k_{t}$-factorization frameworks. Currently, there exists two libraries, such as LHAPDF and TMDLib, to obtain these physical objects. However, there are limitations in both libraries, especially for TMDs, such as restricted customization and extensibility. Users are limited to the implementations provided by these libraries and cannot easily support unconventional PDFs. Moreover, no standard TMD library currently provides a consistent framework for QCD coupling evaluation or for studying uncertainties at the distribution level—features that are important for diagnostic and comparative analyses in phenomenological research.
@@ I Introduction
For TMDLib, these issues manifest differently. Although common metadata (info) files are available for all sets, there is no widely adopted standard for TMD sets regarding the grid or file structure (and, in practice, the grid layout conventions). Interpolation is performed according to the prescriptions defined by the respective set authors, as described in the TMDLib manual. The library is, in fact, a collection of independent components, each handling its specific TMD set under a common Abstract Programming Interface (API). Due to this lack of a standardized grid/file procedure, adding support for a new TMD set remains a non-trivial task. Additionally, while uncertainty members (where provided) are stored in a standardized way, uncertainty propagation cannot be performed within the library itself at the level of the distributions and is only feasible through TMDplotter [16]. This tool enables diagnostic analyses of PDF- or TMD-level uncertainties in a manner consistent with LHAPDF. In PDFxTMDLib, this functionality is extended by introducing an explicit ErrorType field in the metadata, following the LHAPDF convention, to enable consistent handling of PDF- and TMD-level uncertainty information.
@@ II.2.1 The PDFSet Class
  • AlphaQCDMu2(mu2), Computes the strong coupling $\alpha_{s}$ at a given scale $\mu^{2}$.
@@ II.5 Grid Data Formats and File Handling
PDFxTMDLib maintains compatibility with lhagrid1 file formats of LHAPDF while introducing a new file format for TMDs which is an extension of the lhagrid1 format to address the lack of standardization in the TMD community. We name the new file format lhagrid_tmd1. This extended format maintains structural similarity to lhagrid1 but accommodates the additional dimensionality of TMDs by organizing data in grids of $x$, $k_{t}^{2}$, $\mu^{2}$, and flavors. For implementation simplicity, the current version utilizes a single subgrid structure, in contrast to the multiple subgrids supported in the lhagrid1 format.
At the present stage, PDFxTMDLib provides validated support for a subset of TMD sets, specifically the parton-branching (PB) family: PB-LO-HERAI+II-2020-set1, PB-LO-HERAI+II-2020-set2, PB-NLO+QED-HERAI+II-set1, PB-NLO+QED-HERAI+II-set2, PB-NLO-HERAI+II-2023-set2-qs=0.74, and PB-NLO-HERAI+II-2023-set2-qs=1.04. For these sets, the exact grid data files have been converted into the lhagrid_tmd1 format using set-specific conversion tools, and validated by direct numerical comparison with TMDLib, as shown in section III and figure 8, where excellent agreement is observed.
At present, PDFxTMDLib does not provide a general mechanism for generating or converting arbitrary TMDLib grids into the lhagrid_tmd1 format. Grid conversion and validation are currently limited to the PB sets listed above, for which the grid structure and interpolation prescriptions are well defined. Support for additional TMD formats will require dedicated, set-specific development and validation.
In addition, PDFxTMDLib maintains partial support for some TMD sets of TMDLib through specialized IReader and IInterpolator implementations for the allflavorUpdf format. However, due to the lack of a standard grid in TMDLib, this support is partial. For these other TMD sets, TMDLib remains the only choice, as it directly uses author-provided implementations that guarantee exact reproduction and correct handling of set-specific theoretical assumptions.
The goal of PDFxTMDLib is therefore not to replace TMDLib, but to provide a unified and extensible framework that enables standardized storage and access to TMD grids where possible, while lowering the technical barrier for adding new TMD sets in the future. The library’s support for TMD sets will expand progressively, with updates published published at www.pdfxtmdlib.org.
@@
One additional improvement offered by PDFxTMDLib is the extension of uncertainty and QCD coupling functionality to the TMD domain. While these features have long been standard in collinear PDF libraries, PDFxTMDLib introduces them to TMD analyses for the first time, enabling diagnostic and comparative studies of uncertainties directly at the distribution level. It should be noted, however, that these uncertainty estimates correspond to variations among the members of a TMD or PDF set and are not intended to represent uncertainties of physical cross sections. For cross-section calculations, all individual set members must be propagated through the full phenomenological computation. In this respect, the implementation in PDFxTMDLib follows the same philosophy as LHAPDF and TMDplotter, where uncertainty information is exposed at the level of the distributions themselves.
@@
In this structure, <nnnn><nnnn> denotes a four-digit member identifier. For instance, ‘CT18NLO_0003.dat’ represents the third member of the ‘CT18NLO’ PDF set. The central metadata for each set is stored in a YAML-formatted info file bearing the set name.
@@ III.1 Drell-Yan Process Simulations with PDFxTMDLib and LHAPDF
To assess the validity of PDFxTMDLib, we compare the total and differential cross sections with those obtained using LHAPDF. The total Drell–Yan cross section is measured to be $1.695\times 10^{-6}$ mb in both libraries, with statistical uncertainties of $1.453\times 10^{-10}$ mb, demonstrating perfect agreement. Differential cross sections, presented as histograms of the dilepton invariant mass, rapidity, and transverse momentum, also show excellent consistency between the two implementations. These comparisons are displayed in Figure 6, where the overlaid distributions confirm the alignment across all kinematic variables. This agreement indicates that PDFxTMDLib provides a reliable alternative to LHAPDF.
It is worth noting, however, that a direct performance comparison with TMDLib is not straightforward, as TMDLib consists of multiple independently developed implementations corresponding to different TMD sets, each exhibiting distinct computational characteristics. Consequently, a uniform benchmarking comparison across its various components is non-trivial. For this reason, we do not attempt such a comparison here, although interested readers may pursue it independently. In the following subsection, we examine the validity of PDFxTMDLib in more detail by comparing both its collinear PDFs (cPDFs) and TMDs against the corresponding results from LHAPDF and TMDLib.
Figure 6: Comparison of differential cross sections for the Drell–Yan process simulated with PYTHIA using PDFxTMDLib (solid lines) and LHAPDF (dashed lines) with the MSHT20nlo_as120 PDF set. The upper panels show the absolute distributions, while the lower panels show the ratio $\mathrm{PDFxTMDLib}/\mathrm{LHAPDF}$; the horizontal line indicates unity. From top left to bottom, the observables shown are the dilepton rapidity $y_{\ell\ell}$, transverse momentum $p_{T,\ell\ell}$, and invariant mass $m_{\ell\ell}$.
@@ IV Installation and Usage Guide
This section provides instructions for building PDFxTMDLib and integrating it into research workflows. The source code is publicly available on GitHub at https://github.com/Raminkord92/PDFxTMD, and the latest versions can also be downloaded from the project webpage at https://pdfxtmdlib.org/downloads/ (e.g. https://pdfxtmdlib.org/downloads/PDFxTMD-main.zip). For any questions regarding the installation or usage of the library, users can contact us by email or via the official GitHub discussions page: https://github.com/Raminkord92/PDFxTMD/discussions.
@@ IV.1.2 Build Process
cmake -S . -B build -DCMAKE_INSTALL_PREFIX=/path/to/install -DCMAKE_BUILD_TYPE=Release
cmake --build build
@@ IV.1.3 Installation
To install the library, execute:
cmake --build build --target install
This installs headers, libraries, and CMake configuration files under the chosen prefix. If no prefix is specified, CMake will use its default install location (which may be a system directory such as /usr/local on Linux/macOS).
@@ IV.4.3 Uncertainty and Correlation Analysis
Important note: The uncertainty values returned by Uncertainty(...) refer only to the PDF or TMD distributions and not to physical cross sections.
@@ V Conclusion
Numerical validation, including Drell-Yan process simulations with PYTHIA and direct comparisons with LHAPDF and TMDLib, demonstrates that PDFxTMDLib delivers consistent and accurate results, alongside modest performance improvements in specific scenarios. The library further enhances phenomenological research by introducing novel features for TMDs, such as uncertainty quantification at distribution level and QCD coupling calculations, which were previously unavailable in a standardized form. Additionally, we also introduce a new lhagrid_tmd1 file format for TMDs which is an extension of lhagrid file format to facilitate and standardize calculations for TMDs.
@@
Data Availability Statement
No Data associated in the manuscript.